Diffusion Morning Report 04/08/11

Ted Berg tberg at stanford.edu
Fri Apr 8 07:20:23 PDT 2011


Good Morning All,
     Lo0oking good this morning LTO has door issues. Epi had some 
mapping issues. SEM 4160 PM is to begin today.Thermco 2 has no Tlc 
flow.No other issues at this time.

					
	Date: 	4/8/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	up 				
Tylan 2 	up 				
Tylan 3 	up 				
Tylan 4 	up 				
Tylan 5-BBr3 	up 				
Tylan 6-POCl3 	up 				
Tylan 7-TEOS2 	up 		.25/30 		
Tylan 8-SiGe 	up 		.88/30 		
Tylan 9-Poly 	up 		.12/30 		
Tylan 10-Nitride 	up 		9.92/30 		
Tylan 11-LTO 	yellow 	Door close-O2 flow issues 	8.38/30 		1
Tylan 12-FGA 	up 				
Thermco 1 	up 				
Thermco 2 	yellow 	no tlc flow 			
Thermco 3 	future 				
Thermco4 	future 				
Thermco Poly1 	up 				
Thermco Poly2 	up 				
Thermco Nitride1 	up 				
Thermco bpsg 	up 				
AMAT Epi 	yellow 	mapping and SiH4 flow issues 			2
TEL Spa 	up 				
Burn Box 	up 				
Ebeam 	down 	cryo compresssor dead 			
ZYGO 	up 	
			
FGA2 	up 				
SiH4 for furnaces 		A on line 469 PSI 	changeover working well		


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110408/ab0fa3f9/attachment.html>


More information about the diffusion mailing list